OpenStudy (devan):

how do i solve : -7x-8y=7-11y

OpenStudy (anonymous):

isolate y. move -11y to the other side to get -7x+3y=7.... Then move the x to the other side to get 3y=7x+7 then divide by 3 y=7x/3 + 7/3

OpenStudy (devan):

so the answer would be 7/3

OpenStudy (anonymous):

7/3 is your slope, and in this case it is your y intercept as well. I don't know what the question is asking so I can't say that that is the answer.

OpenStudy (devan):

the question asks find the slope of -7x-8y=7-11y